OpenStudy (iwanttogotostanford):

NEED HELP W/ THIS ACT Q. PLEASE: If m and n are positive integers such that 2/3 of m is equal to 1/2 of n, then m is what fractional part of n?? A. 1/3 B. 3/5 C. 3/4 D. 4/3 E. 5/3

OpenStudy (retireed):

m and n (2/3)m = (1/2)n m = (3/2)(1/2) n m = (3/4)n I down to C or D. I would say m is _______ of n.

OpenStudy (mathmale):

Form the ratio \[\frac{ m }{ n }=\frac{ \frac{ 3n }{ 4 } }{ n }\] and reduce this.

OpenStudy (iwanttogotostanford):

@mathmale do I choose numbers for m and n and then solve it?

OpenStudy (mathmale):

Try cancelling those two n's. The whole point here was to eliminate m, so no, do not choose a valule for m.